OpenStudy (experimentx):

Evaluate:: \[ \huge \lim_{k \rightarrow \infty} \sum_{n = k+1}^{2k} {1 \over n}\]

OpenStudy (anonymous):

ln 2

OpenStudy (australopithecus):

http://www.wolframalpha.com/input/?i=lim+k-%3E+infinity+summation+n%3D1%2Bk+to+2k+of+1%2Fn no you are not mukushla at least according to wolfram alpha

OpenStudy (anonymous):

ln(2) ! yes !

OpenStudy (australopithecus):

how did you get that just wondering this is too advanced for me probably so I don't know why I bother asking

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\frac{1}{x+1}dx\] @mukushla

OpenStudy (anonymous):

sorry i was right first time...

OpenStudy (experimentx):

log 2

OpenStudy (unklerhaukus):

would the answer be a function of \(k\)?

OpenStudy (experimentx):

this is pretty unintuitive.

OpenStudy (experimentx):

k -> inf

OpenStudy (unklerhaukus):

i see the limit now,

OpenStudy (experimentx):

still ... this is pretty unintuitive... atleast for me when i saw it for the first time.

OpenStudy (unklerhaukus):

can we break the sum into two sums?

OpenStudy (anonymous):

\[ \lim_{k \rightarrow \infty} \sum_{n = k+1}^{2k} {1 \over n}= \lim_{k \rightarrow \infty} (\frac{1}{1+k}+\frac{1}{2+k}+...+\frac{1}{2k})\\=\lim_{k \rightarrow \infty} \frac{1}{k}(\frac{1}{1+\frac{1}{k}}+\frac{1}{1+\frac{2}{k}}+...+\frac{1}{1+\frac{k}{k}})=\int\limits_{0}^{1} \frac{1}{1+x} dx=\ln 2 \]

OpenStudy (experimentx):

yep ... i believe we can. but this method can be quite difficult ...

OpenStudy (unklerhaukus):

\[\sum_{n=k+1}^{2k}=\sum_{n=0}^{2k}-\sum_{n=0}^{k+1}\]

OpenStudy (experimentx):

this is related to http://en.wikipedia.org/wiki/Euler%E2%80%93Mascheroni_constant

OpenStudy (experimentx):

*this method the other is a lot simpler.

OpenStudy (anonymous):

or http://en.wikipedia.org/wiki/Riemann_integral

OpenStudy (experimentx):

Riemann sums is easier also \[ \sum_{n = k+1}^{2k}{1 \over k{ n \over k}} = \int_1^2 {}1 {\over x } dx\]

OpenStudy (experimentx):

*1/x dx

OpenStudy (unklerhaukus):

hmm, i dont get it

OpenStudy (experimentx):